OpenStudy (anonymous):

the product of two integers m and k is 12. if each integer is increased by 1 what could be the new product?

OpenStudy (dumbcow):

mk = 12 (m+1)(k+1) = mk + m +k+1 = 13+m+k

OpenStudy (dumbcow):

m,k could be (1,12) (2,6) or (3,4)

OpenStudy (anonymous):

@dumbcow u got it!

OpenStudy (anonymous):

m,k can be 1,12 2,6 3,4 so increased by one would be 2x13=26 3x7=21 4x5=20